OpenStudy (anonymous):

difference of 9√7 and √28?

OpenStudy (mertsj):

\[\sqrt{28}=\sqrt{4(7)}=2\sqrt{7}\]

OpenStudy (mertsj):

Does that help?

OpenStudy (anonymous):

ummm not really?

OpenStudy (mertsj):

So now the problem is : \[9\sqrt{7}-2\sqrt{7}\]

OpenStudy (anonymous):

is it 11 square root of 7?

OpenStudy (mertsj):

Does that help?

OpenStudy (mertsj):

You subtract to find the difference between two numbers.

OpenStudy (anonymous):

oh so square root of 7?

OpenStudy (anonymous):

i mean 7 square root of 7?

OpenStudy (mertsj):

That's better.